OpenStudy (anonymous):

if x varies directly as y, and x=9 when y=2, find the x when y is 6

OpenStudy (anonymous):

if x varies directly as y then we can say x=ky where k is consatant of proportionality

OpenStudy (anonymous):

hence we can say 9=k*2 so k= 9/2 now when y=6 x = ky x= (9/2) *6 x=27